# DRPO: Efficient Reasoning via Decoupled Reward Policy Optimization

### v1

@@ 1 Introduction
However, existing LRMs suffer from overthinking with lengthy and redundant reasoning paths. As demonstrated by Chen et al. (2024), reasoning models like DeepSeek-R1 (Guo et al., 2025) need to generate about 1,000 tokens to answer “what is the answer of 2 plus 3”, while only around 10 tokens are needed for non-reasoning models. Such overly generated reasoning paths raise significant issues, leading to substantially increased computational cost and longer inference time. Numerous studies have been conducted to explore ways to eliminate redundant reasoning and improve the reasoning efficiency. A popular strategy is to introduce explicit reward shaping with length penalties in RL to guide the model toward concise reasoning (Arora and Zanette, ; Huang et al., ; Xiang et al., ; Aggarwal and Welleck, ), e.g., penalizing rewards of correct answers based on reasoning length to encourage shorter reasoning. Nevertheless, almost all existing methods fall short in preserving performance while shortening reasoning, causing substantial performance loss. This raises a key question: how to guide RL for efficient reasoning with minimal performance drop?
@@ 2 Related work
The breakthrough came with DeepSeek-R1, which revealed that large reasoning models (LRMs) trained via large-scale reinforcement learning, particularly GRPO (Shao et al., 2024), can autonomously acquire advanced reasoning behaviors such as branching, verification, and backtracking. This success inspired numerous follow-up studies: some aimed at reproducing the effectiveness of GRPO (Wen et al., 2025; Luo et al., 2025b; He et al., 2025), while others investigated its limitations and proposed refinements to further enhance reasoning performance (Liu et al., 2025c; Yu et al., 2025; Li et al., 2025a). In parallel, many open-weight reasoning models have adopted GRPO, including Qwen-3 (Yang et al., 2025a), GLM-4.5 (Team et al., 2025a), K2-Think (Cheng et al., 2025), and Goedel-solver (Lin et al., 2025), among others. Despite these advances, existing LRMs often suffer from overthinking—producing unnecessarily long and redundant reasoning even for simple problems. This work aims to directly address this issue by training LRMs to reason both efficiently and effectively.
@@
Efficient Reasoning in LRMs. To address the issue of overthinking in LRMs, a variety of techniques have been proposed (Sui et al., 2025; Yue et al., 2025), including (1) training-free methods, which shorten the reasoning paths via prompt (Aytes et al., 2025; Han et al., 2024) or manipulating the decoding process (Yang et al., 2025b; Yong et al., 2025; Wang et al., 2025a; Liu et al., 2025a; Wang et al., 2025b); (2) Supervised Fine-tuning (SFT) methods, which rely on compressed reasoning datasets for finetuning. These datasets are curated via token-level selection (Yuan et al., 2025; Xia et al., 2025; Zhuang et al., 2025), step-level selection (Xiao et al., 2025; Cui et al., 2025; Wang et al., 2025c), path-level selection (Munkhbat et al., 2025; Ghosal et al., 2025); (3) RL-based methods, which carefully design reward mechanisms to guide the model to reason efficiently (Hou et al., ; Aggarwal and Welleck, ; Arora and Zanette, ; Xiang et al., ; Luo et al., ; Huang et al., ; Yi et al., 2025; Liu et al., 2025b; Xu et al., 2025; Fang et al., 2025; Li et al., 2025b).
@@
Among these techniques, RL-based methods have been demonstrated to be one of the most effective approaches due to their scalability and flexibility. Specifically, L1 (Aggarwal and Welleck, ) integrates a length constraint into the reward function to encourage higher performance while meeting the length goal specified in the prompt. Arora and Zanette employs online RL with a length penalty based on the distribution of correct answers, penalizing correct responses longer than the average while encouraging those shorter than the average. ALP (Xiang et al., ) adaptively adjusts length penalties according to problem difficulty, measured by pass rate, assigning stronger penalties to high pass-rate problems to discourage overlong reasoning. HAPO (Huang et al., ) keeps track of the minimum length of correct responses for each question, penalizing outputs that exceed this length while rewarding those that are shorter. Nevertheless, these methods all suffer from misleading learning signals and fall short in preserving performance while shortening reasoning, due to the limitation of their adopted relative advantage function.
@@ 3 Limitation of incorporating length penalty into GRPO
Let us consider incorporating a length penalty into GRPO. Existing RL methods with length control reveal a shared principle: penalizing rewards of correct answers based on reasoning length to encourage shorter reasoning, e.g., $r(o|q)=r_{c}(o|q)-r_{l}(o|q)$, where $r_{l}(\cdot)$ is a length-based cost or reward function (refer to Table 2 for detailed formulations). The consequence is that the reward of a correct with long output will be shifted down relatively. Incorporating this reward into GRPO’s group relative advantage calculation may bias the intended effect, misleading the learning process. Let us consider an illustrative example in Figure 1. Suppose there are six generated outputs with different lengths, whose correctness rewards are $[1,1,1,0,0,0]$ and corresponding lengths are $[2000,2500,4000,2800,3800,3200]$. After combining the length reward with correctness reward, the combined reward of each answer becomes $[0.73,0.6,0.2,0,0,0]$  [^1] These values are calculated with the formula proposed in  (Arora and Zanette, ). We note that various reward combination designs in the literature lead to the same issue. Refer to Appendix A.3 for a detailed discussion. . So far, it looks like that all the designs work well since short correct answers have larger positive rewards and longer correct answers have smaller positive rewards while incorrect answers have zero rewards. However, when computing group relative advantage $A(o|q)$ in GRPO (i.e., Eqn. (2)), the advantage for the third correct response shifts from 1 (without length penalty) to -0.17 (with length penalty added). This negative signal will discourage valid reasoning and create a significant optimization barrier, which may substantially harm performance.
@@ 4.2 Decoupled Reward Policy Optimization
It is notable that the final objective only relies on the on-policy data, and it has an explanation that each positive data is assigned with a weight $\omega(o|q)=\frac{\exp(r_{l}(o)/\lambda)}{\mathbb{E}_{o\sim\pi_{\text{old}}^{+}(\cdot|q)}\exp(r_{l}(o)/\lambda)}$ informed by its length but normalized only withing the positive data. It is notable that when $\lambda=+\infty$, then $\omega(o|q)=1$ and the above objective reduces to that of DisCO in (3).
@@ 5 Experiments
Datasets. We validate our method on mathematical reasoning tasks. Specifically, we train models on the DeepScaleR-Preview-Dataset (Luo et al., 2025c), which consists of approximately 40.3k question-answer pairs sourced from AIME problems from 1984 to 2023, AMC problems before 2023, Omni-MATH (Gao et al., 2024) and Still (Min et al., 2024) datasets. We evaluate all the models on math problems with different levels of difficulty, including (a) easy level: GSM8K (Cobbe et al., 2021), (b) medium level: MATH-500 (Hendrycks et al., 2021), (c) hard level: OlympiadBench (He et al., 2024), and (d) very hard level: AIME (aggregating 2024 and 2025).
@@
Models. We adopt two reasoning models as our base models: DeepSeek-R1-Distill-Qwen-1.5B model and DeepSeek-R1-Distill-Qwen-7B, and conduct RL fine-tuning from them.
@@
Baselines. We compare our methods with six of the most recent state-of-the-art efficient reasoning methods, including (1) the method in (Arora and Zanette, ), which integrates length reshaped rewards into the RLOO advantage function, referred to as RLOO-LP; (2) ALP (Xiang et al., ), which uses a length penalty in GRPO that is the length scaled by the solving rate of each question; (3) HAPO (Huang et al., ), which penalizes the responses longer than the shortest correct answer in the history while rewarding those that are shorter; (4) L1-max (Aggarwal and Welleck, ), which is a reasoning language model that produces outputs satisfying a maximum length constraint given in the prompt; (5) ShorterBetter (SB) (Yi et al., 2025), which aims to match Sample Optimal Length defined as the shortest correct response among multiple generations; (6) LASER-D (Liu et al., 2025b), which employs a step length reward function based on difficulty-aware dynamic target length. We train models for methods (1)-(3) using the experimental settings described below. For methods (4)-(6), we evaluate the models provided in their original works. All the compared models were finetuned from the same base models on the DeepScaleR-Preview dataset, except L1-max, which was trained on DeepScaleR-1.5B-Preview. We summarize different reward designs of the above baselines in Table 2 in Appendix.
@@
Training Details. For all the training, we employ the AdamW optimizer with a weight decay of 0.01 and set the learning rate to a constant $2\mathrm{e}^{-6}$ for 1.5B model and $1\mathrm{e}^{-6}$ for 7B model, following  Li et al. (2025a). We set the batch size to 128 for each step of RL, the mini-batch size to 32 for each iteration of model update, and sample 8 responses per question for training. For RLOO-LP, we tune their weight parameter $\alpha\in\{0.05,0.1,0.2\}$. For ALP, we tune their penalty weight $\beta\in\{1\mathrm{e}^{-9},1\mathrm{e}^{-8},1\mathrm{e}^{-7}\}$. For HAPO, we tune their weight parameter $w\in\{0.01,0.1,1\}$. For the proposed method, we tune $\lambda\in\{0.5,0.2,0.1\}$. These parameters serve the same role that controls the tradeoff between efficiency and accuracy. For all other hyperparameters, we follow the default values from their official papers. Details are provided in Appendix A.2. The generation budget is limited to 8k tokens for both training and evaluation.
@@
where $\alpha,\beta,\gamma>0$, $\Delta_{\text{Length}}=\frac{\text{Length}_{\text{ref}}-\text{Length}_{\text{model}}}{\text{Length}_{\text{ref}}}$ and $\Delta_{\text{Acc}}=\frac{\text{Acc}_{\text{model}}-\text{Acc}_{\text{ref}}}{\text{Acc}_{\text{ref}}}$, where the quantities with subscript “ref" means each method’s baseline that does not consider length reward, and that with subscript “model" means the model for evaluation. In our experiments, we use the default values $\alpha=1,\beta=3$ same as (Luo et al., 2025a), but set $\gamma=10$ to emphasize the importance of minimizing performance degradation.
@@ 5.1 Visualization of Learning Process
Figure 3: Comparison of performance-efficiency trade-off. Left is for fine-tuning 1.5B model and right is for fine-tuning 7B model. Grey lines represent the base model performance before finetuning, with generation length of 4698 for 1.5B model and 4119 for 7B model. Squares denote models trained with reference methods without length penalties (i.e., $\lambda$$\infty$ for DRPO, $\alpha=0$ for RLOO-LP, $\beta=0$ for ALP, $w=0$ for HAPO). Triangles denote the models trained by other works.
@@ 5.2 Comparison with baselines
Trade-off between performance and efficiency. We present the trade-off between performance and efficiency for various methods in Figure 3, where the averaged pass@1 over four math datasets of different difficulty is reported. We observe that our proposed DRPO consistently achieves significantly higher performance than all baselines on finetuning 1.5B and 7B models, including the models trained by other work. Notably, relative to the reference learning method (square marker), the proposed DRPO on finetuning 7B model in Figure 3 (Right) efficiently reduces reasoning length from 3053 to 1502 (51% length reduction) with 2.6% loss of performance via varying $\lambda$ to control the tradeoff, demonstrating the effectiveness of DRPO to reduce reasoning length while preserving the reasoning capability. In contrast, all the efficient reasoning baselines suffer from severe performance degradation when the reasoning length is reduced. For example, RLOO-LP reduces reasoning length from 2975 to 1841 (38% length reduction) but incurs a 7.1% loss in performance on finetuning 7B model, with ALP showing a similar trend. Compared with DRPO, these methods trade off more performance for less reduction in length, highlighting the superiority of our method.
Evaluating Trade-off via AES. Additionally, we directly quantify the effectiveness of our proposed method in enhancing reasoning efficiency with minimal performance drop, using the Accuracy-Efficiency Score (AES). AES is positive when the model reduces output length while maintaining or enhancing accuracy, and negative when accuracy deteriorates. For fair comparison, reference model in AES is each method’s counterpart without length reward, i.e., $\lambda$$\infty$ for DRPO, $\alpha=0$ for RLOO-LP, $\beta=0$ for ALP, $w=0$ for HAPO. We present the best AES score for each method in Table 1 and defer detailed results in Appendix A.4. From Table 1, we observe that all the baseline methods exhibit a negative AES score for both 1.5B and 7B models, indicating the inefficiency of existing methods in reducing reasoning length while preserving performance. In contrast, DRPO consistently achieves a positive AES score for finetuning both 1.5B and 7B models, highlighting its capability of improving reasoning efficiency while maintaining performance.
@@
(a) Performance-efficiency tradeoff on individual datasets with increasing difficulty levels from left to right. (a) is for finetuning 1.5B model and (b) is for finetuning 7B model. Squares denote models trained with reference methods without length penalties (i.e., $\lambda$$\infty$ for DRPO, $\alpha=0$ for RLOO-LP, $\beta=0$ for ALP, $w=0$ for HAPO).

### v2

@@ 1 Introduction
However, existing LRMs suffer from overthinking with lengthy and redundant reasoning paths. As demonstrated by Chen et al. (2024), reasoning models like DeepSeek-R1 (Guo et al., 2025) need to generate about 1,000 tokens to answer “what is the answer of 2 plus 3”, while only around 10 tokens are needed for non-reasoning models. Such overly generated reasoning paths raise significant issues, leading to substantially increased computational cost and longer inference time. Numerous studies have been conducted to explore ways to eliminate redundant reasoning and improve the reasoning efficiency. A popular strategy is to introduce explicit reward shaping with length penalties in RL to guide the model toward concise reasoning (Arora and Zanette, 2025; Huang et al., 2025; Xiang et al., 2025; Aggarwal and Welleck, 2025), e.g., penalizing rewards of correct answers based on reasoning length to encourage shorter reasoning. Nevertheless, almost all existing methods fall short in preserving performance while shortening reasoning, causing substantial performance loss. This raises a key question: how to guide RL for efficient reasoning with minimal performance drop?
@@ 2 Related work
The breakthrough came with DeepSeek-R1, which revealed that large reasoning models (LRMs) trained via large-scale reinforcement learning, particularly GRPO (Shao et al., 2024), can autonomously acquire advanced reasoning behaviors such as branching, verification, and backtracking. This success inspired numerous follow-up studies: some aimed at reproducing the effectiveness of GRPO (Wen et al., 2025; Luo et al., 2025d; He et al., 2025), while others investigated its limitations and proposed refinements to further enhance reasoning performance (Yu et al., 2025; Li et al., 2025a; Chen et al., 2025; Zheng et al., 2025). In parallel, many open-weight reasoning models have adopted GRPO, including Qwen-3 (Yang et al., 2025a), GLM-4.5 (Team et al., 2025a), K2-Think (Cheng et al., 2025), and Goedel-solver (Lin et al., 2025), among others. Despite these advances, existing LRMs often suffer from overthinking—producing unnecessarily long and redundant reasoning even for simple problems. This work aims to directly address this issue by training LRMs to reason both efficiently and effectively.
@@
Efficient Reasoning in LRMs. To address the issue of overthinking in LRMs, a variety of techniques have been proposed (Sui et al., 2025; Yue et al., 2025), including (1) training-free methods, which shorten the reasoning paths via prompt (Aytes et al., 2025; Han et al., 2024) or manipulating the decoding process (Yang et al., 2025b; Yong et al., 2025; Wang et al., 2025a; Liu et al., 2025a; Wang et al., 2025c); (2) Supervised Fine-tuning (SFT) methods, which rely on compressed reasoning datasets for finetuning. These datasets are curated via token-level selection (Yuan et al., 2025; Xia et al., 2025; Zhuang et al., 2025), step-level selection (Xiao et al., 2025; Cui et al., 2025; Wang et al., 2025b), path-level selection (Munkhbat et al., 2025; Ghosal et al., 2025); (3) RL-based methods, which carefully design reward mechanisms to guide the model to reason efficiently (Hou et al., 2025; Aggarwal and Welleck, 2025; Arora and Zanette, 2025; Xiang et al., 2025; Luo et al., 2025b; Huang et al., 2025; Yi et al., 2025; Liu et al., 2025b; Xu et al., 2025; Fang et al., 2025; Li et al., 2025b).
@@
Among these techniques, RL-based methods have been demonstrated to be one of the most effective approaches due to their scalability and flexibility. Specifically, L1 (Aggarwal and Welleck, 2025) integrates a length constraint into the reward function to encourage higher performance while meeting the length goal specified in the prompt. Arora and Zanette (2025) employs online RL with a length penalty based on the distribution of correct answers, penalizing correct responses longer than the average while encouraging those shorter than the average. ALP (Xiang et al., 2025) adaptively adjusts length penalties according to problem difficulty, measured by pass rate, assigning stronger penalties to high pass-rate problems to discourage overlong reasoning. HAPO (Huang et al., 2025) keeps track of the minimum length of correct responses for each question, penalizing outputs that exceed this length while rewarding those that are shorter. Nevertheless, these methods all suffer from misleading learning signals and fall short in preserving performance while shortening reasoning, due to the limitation of their adopted relative advantage function.
Discriminative Learning for LRMs. Discriminative learning is a classical paradigm applied widely to traditional tasks like classifications (Bishop and Nasrabadi, 2006; Yang and Ying, 2022) and rankings (Burges et al., 2005; Cao et al., 2007). These methods follow the principle of raising scores for positive (correct) samples while lowering scores for negative (incorrect) ones. Recently, several works have explored applying the principle of discriminative learning to LRM training. For example,  Li et al. (2025a) proposed a discriminative constrained RL framework with verifiable binary rewards to finetune LRMs. Lyu et al. (2025) and Bai et al. (2025) utilize discriminative loss for behavior cloning on positive samples and policy gradient on negative samples. Su et al. (2025) employs a discriminative learning approach based on positive-negative pairs for reasoning tasks. However, these methods are limited to binary accuracy rewards and don’t address the challenge of overthinking in LRMs.
@@ 3 Limitation of incorporating length penalty into GRPO
Let us consider incorporating a length penalty into GRPO. Existing RL methods with length control reveal a shared principle: penalizing rewards of correct answers based on reasoning length to encourage shorter reasoning, e.g., $r(o|q)=r_{c}(o|q)-r_{l}(o|q)$, where $r_{l}(\cdot)$ is a length-based cost or reward function (refer to Table 2 for detailed formulations). The consequence is that the reward of a correct with long output will be shifted down relatively. Incorporating this reward into GRPO’s group relative advantage calculation may bias the intended effect, misleading the learning process. Let us consider an illustrative example in Figure 1. Suppose there are six generated outputs with different lengths, whose correctness rewards are $[1,1,1,0,0,0]$ and corresponding lengths are $[2000,2500,4000,2800,3800,3200]$. After combining the length reward with correctness reward, the combined reward of each answer becomes $[0.73,0.6,0.2,0,0,0]$  [^1] These values are calculated with the formula proposed in  (Arora and Zanette, 2025). We note that various reward combination designs in the literature lead to the same issue. Refer to Appendix A.3 for a detailed discussion. . So far, it looks like that all the designs work well since short correct answers have larger positive rewards and longer correct answers have smaller positive rewards while incorrect answers have zero rewards. However, when computing group relative advantage $A(o|q)$ in GRPO (i.e., Eqn. (2)), the advantage for the third correct response shifts from 1 (without length penalty) to -0.17 (with length penalty added). This negative signal will discourage valid reasoning and create a significant optimization barrier, which may substantially harm performance.
@@ 4.2 Decoupled Reward Policy Optimization
It is notable that the final objective only relies on the on-policy data, and it has an explanation that each positive data is assigned with a weight $\omega(o|q)=\frac{\exp(r_{l}(o)/\lambda)}{\mathbb{E}_{o\sim\pi_{\text{old}}^{+}(\cdot|q)}\exp(r_{l}(o)/\lambda)}$ informed by its length but normalized only within the positive data. It is notable that when $\lambda=+\infty$, then $\omega(o|q)=1$ and the above objective reduces to that of DisCO in (3).
@@ 5 Experiments
Datasets. We validate our method on mathematical reasoning tasks. Specifically, we train models on the DeepScaleR-Preview-Dataset (Luo et al., 2025c), which consists of approximately 40.3k question-answer pairs sourced from AIME problems from 1984 to 2023, AMC problems before 2023, Omni-MATH (Gao et al., 2024) and Still (Min et al., 2024) datasets. We evaluate all the models on math problems with different levels of difficulty, including (a) easy level: GSM8K (Cobbe et al., 2021), (b) medium level: MATH-500 (Hendrycks et al., 2021), (c) hard level: OlympiadBench (He et al., 2024), and (d) very hard level: AIME (aggregating 2024 and 2025). To verify the generalizability of our method to non-mathematical reasoning tasks, we also conducted experiments on K&K logic puzzle dataset (Xie et al., 2024), which is included in Appendix A.4.
@@
Models. We adopt three reasoning models as our base models: DeepSeek-R1-Distill-Qwen-1.5B model, DeepSeek-R1-Distill-Qwen-7B, and DeepSeek-R1-Distill-Llama-8B, and conduct RL fine-tuning from them.
@@
Baselines. We compare our methods with six of the most recent state-of-the-art efficient reasoning methods, including (1) the method in (Arora and Zanette, 2025), which integrates length reshaped rewards into the RLOO advantage function, referred to as RLOO-LP; (2) ALP (Xiang et al., 2025), which uses a length penalty in GRPO that is the length scaled by the solving rate of each question; (3) HAPO (Huang et al., 2025), which penalizes the responses longer than the shortest correct answer in the history while rewarding those that are shorter; (4) L1-max (Aggarwal and Welleck, 2025), which is a reasoning language model that produces outputs satisfying a maximum length constraint given in the prompt; (5) ShorterBetter (SB) (Yi et al., 2025), which aims to match Sample Optimal Length defined as the shortest correct response among multiple generations; (6) LASER-D (Liu et al., 2025b), which employs a step length reward function based on difficulty-aware dynamic target length. We train models for methods (1)-(3) using the experimental settings described below. For methods (4)-(6), we evaluate the models provided in their original works. All the compared models were finetuned from the same base models on the DeepScaleR-Preview dataset, except L1-max, which was trained on DeepScaleR-1.5B-Preview. We summarize different reward designs of the above baselines in Table 2 in Appendix.
@@
Training Details. For all the training, we employ the AdamW optimizer with a weight decay of 0.01 and set the learning rate to a constant $2\mathrm{e}^{-6}$ for 1.5B model, $1\mathrm{e}^{-6}$ for 7B model, and $5\mathrm{e}^{-7}$ for 8B model, following  Li et al. (2025a). We set the batch size to 128 for each step of RL, the mini-batch size to 32 for each iteration of model update, and sample 8 responses per question for training. For RLOO-LP, we tune their weight parameter $\alpha\in\{0.05,0.1,0.2\}$. For ALP, we tune their penalty weight $\beta\in\{1\mathrm{e}^{-9},1\mathrm{e}^{-8},1\mathrm{e}^{-7}\}$. For HAPO, we tune their weight parameter $w\in\{0.01,0.1,1\}$. For the proposed method, we tune $\lambda\in\{0.5,0.2,0.1\}$. These parameters serve the same role that controls the tradeoff between efficiency and accuracy. For all other hyperparameters, we follow the default values from their official papers. Details are provided in Appendix A.2. The generation budget is limited to 8k tokens for both training and evaluation.
@@
where $\alpha,\beta,\gamma>0$, $\Delta_{\text{Length}}=\frac{\text{Length}_{\text{ref}}-\text{Length}_{\text{model}}}{\text{Length}_{\text{ref}}}$ and $\Delta_{\text{Acc}}=\frac{\text{Acc}_{\text{model}}-\text{Acc}_{\text{ref}}}{\text{Acc}_{\text{ref}}}$, where the quantities with subscript “ref” means each method’s baseline that does not consider length reward, and that with subscript “model” means the model for evaluation. In our experiments, we use the default values $\alpha=1,\beta=3$ same as (Luo et al., 2025a), but set $\gamma=10$ to emphasize the importance of minimizing performance degradation.
@@ 5.1 Visualization of Learning Process
Figure 3: Comparison of performance-efficiency trade-off. Left is for fine-tuning 1.5B model, middle is for fine-tuning 7B model and right is for fine-tuning 8B model. Grey lines represent the base model performance before finetuning, with generation length of 4698 for 1.5B model, 4119 for 7B model, and 4325 for 8B model. Squares denote models trained with reference methods without length penalties, i.e., $\lambda$$\infty$ (corresponding to DisCO) for DRPO, $\alpha=0$ for RLOO-LP, $\beta=0$ (corresponding to GRPO) for ALP, $w=0$ for HAPO. Triangle, star, and rhombus markers represent models trained by other works.
@@ 5.2 Comparison with baselines
Trade-off between performance and efficiency. We present the trade-off between performance and efficiency for various methods in Figure 3, where the averaged pass@1 over four math datasets of different difficulty is reported. We observe that our proposed DRPO consistently achieves significantly better performance-efficiency trade-off than all baselines on finetuning 1.5B, 7B and 8B models, including the models trained by other work. Notably, relative to the reference learning method (square marker), the proposed DRPO on finetuning 7B model in Figure 3 (Middle) efficiently reduces reasoning length from 3053 to 1502 (51% length reduction) with only 2.6% loss of performance via varying $\lambda$ to control the tradeoff, demonstrating the effectiveness of DRPO to reduce reasoning length while preserving the reasoning capability. In contrast, all the efficient reasoning baselines suffer from severe performance degradation when the reasoning length is reduced. For example, RLOO-LP reduces reasoning length from 2975 to 1841 (38% length reduction) but incurs a 7.1% loss in performance on finetuning 7B model, with ALP showing a similar trend. Compared with DRPO, these methods trade off more performance for less reduction in length, highlighting the superiority of our method.
Evaluating Trade-off via AES. Additionally, we directly quantify the effectiveness of our proposed method in enhancing reasoning efficiency with minimal performance drop, using the Accuracy-Efficiency Score (AES). AES is positive when the model reduces output length while maintaining or enhancing accuracy, and negative when accuracy deteriorates. For fair comparison, reference model in AES is each method’s counterpart without length reward, i.e., $\lambda$$\infty$ for DRPO, $\alpha=0$ for RLOO-LP, $\beta=0$ for ALP, $w=0$ for HAPO. We present the best AES score for each method in Table 1 and defer detailed results in Appendix A.6. From Table 1, we observe that almost all the baseline methods exhibit negative AES scores across 1.5B, 7B, and 8B models, indicating the inefficiency of existing methods in reducing reasoning length while preserving performance. In contrast, DRPO consistently achieves a positive AES score for finetuning 1.5B, 7B, and 8B models, highlighting its capability of improving reasoning efficiency while maintaining performance.
@@
8B Model RLOO-LP 0.619 2249 0.251
ALP 0.637 2358 -0.01
HAPO 0.595 1981 -0.366
DRPO 0.690 1758 0.297
(a) Performance-efficiency tradeoff on individual datasets with increasing difficulty levels from left to right. (a) is for finetuning 1.5B model, (b) is for finetuning 7B model, (c) is for finetuning 8B model. Squares denote models trained with reference methods without length penalties (i.e., $\lambda$$\infty$ (corresponding to DisCO) for DRPO, $\alpha=0$ for RLOO-LP, $\beta=0$ (corresponding to GRPO) for ALP, $w=0$ for HAPO).